17. An exercise physiologist records the mile times in seconds for 10 college distance runners before and after a twelve-week training cycle. Is there evidence at the .05 level that the runners' mile times improved (were lower...) following the twelve weeks' training?

Before After
249.9 250.2
278.8 266.2
254.2 244.7
261.8 252.7
238.8 237.6
267.2 269.7
267.8 256.8
259.8 263.1
254.5 252.1
255.4 243.5

What's the value of the test statistic (based on the After-Before differences)?

18. What's the p-value?

19. Which of the following is(are) the correct critical value(s)?

-1.833

-1.96, 1.96

-2.33

-2.821

-2.262, 2.262

20.

What is the correct decision?

Reject H0

Don't Reject HA

Reject HA

Don't Reject H0
Answer #1

Solution:

(Part 17)

Here, we have to use the paired t test for the population mean difference. The null and alternative hypotheses for this test are given as below:

Null hypothesis: H0: The runners’ average mile times are not improved.

Alternative hypothesis: Ha: The runners’ average mile times are improved.

H0: µd = 0 versus Ha: µd < 0

This is a lower tailed test.

We are given level of significance = α = 0.05

The test statistic formula for this test is given as below:

t = (Dbar - µd)/[Sd/sqrt(n)]

From given data, we have

Dbar = -5.1600

Sd = 6.2593

n = 10

df = n – 1 = 9

α = 0.05

t = (Dbar - µd)/[Sd/sqrt(n)]

t = (-5.1600 - 0)/[ 6.2593/sqrt(10)]

t = -5.16/ 1.9793

t = -2.6069

(Part 18)

P-value = 0.0142

(by using t-table)

(Part 19)

Critical value = -1.833

(by using t-table)

(Part 20)

P-value < α = 0.05

So, we reject the null hypothesis

Reject H0

There is sufficient evidence to conclude that the runners’ average mile times are improved.